OpenStudy (anonymous):

The concentration of a chemical solution is often described as the number of moles of compound per litre of total solution (mol/L). If 11 L of a 1.4 mol/L hydrochloric acid (HCl) solution is mixed with 3 L of a 0.75 mol/L HCl solution, How many moles of HCl are in the final mixture? Round to 2 significant digits.

OpenStudy (anonymous):

I am only guessing here because I haven't done anything exactly like this .. but this is how I would try doing it ...........\[\frac{ 1.4mol \times 11L }{L } + \frac{ 0.75mol \times 3L }{ L } =20.95mol \] rounded to 2 sig figs should be 21mol... I'm not sure if this is correct but it makes sense.